OpenStudy (moonlitfate):

How long to the nearest second does it take a car to accelerate from rest at 3 m/s^2 to cover a distance of 119 meters?

OpenStudy (abhisar):

u=0 s=119 a=3 Plug in the values into \(\sf s=ut+1/2at^2\) and find the value of t

OpenStudy (moonlitfate):

@Abhisar -- Hmm, I don't think I got the right thing.

OpenStudy (moonlitfate):

Wow, defintely didn't get that..

OpenStudy (abhisar):

u got how i did the above calculation ?

OpenStudy (moonlitfate):

You have to isolate t. Then again my math skills could just be really rusty..

OpenStudy (abhisar):

ok i am elaborating it...it's easy \(\sf \because~s=ut+1/2at^2\\ \Rightarrow 119=0\times t+0.5\times 3\times t^2\\ 119=1.5\times t^2\\ t^2=119/1.5\\t^2=79.33\\t=\sqrt{79.33}=8.9s\)

OpenStudy (abhisar):

My first answer was wrong !

OpenStudy (moonlitfate):

Ooooh, I was putting the at^2 in the denominator of the fraction.

OpenStudy (abhisar):

Is it clear now ?

OpenStudy (moonlitfate):

I got 8.906, and since it's the nearest second, it would be 9?

OpenStudy (abhisar):

yes

OpenStudy (moonlitfate):

:) That's so much easier than what my book has.

OpenStudy (abhisar):

:)

OpenStudy (moonlitfate):

I'll be posting more questions. I need th epractice, lol.

OpenStudy (abhisar):

ok :)